£(x - 50) =60,£(y-20) = -20,£(x -50)^2 = 450,£(y - 20)^2 = 200,£(x - 50)(y - 20) = -100


1
Expert's answer
2021-06-23T17:30:42-0400

Solution:

£(x - 50) =60,

"\Rightarrow x=60+50=110"

£(y-20) = -20,

"\Rightarrow y=-20+20=0"

£(x -50)^2 = 450,

"\Rightarrow x=\pm\sqrt{450}+50=\pm15\sqrt{2}+50"

£(y - 20)^2 = 200,

"\Rightarrow y=\pm\sqrt{200}+20=\pm10\sqrt{2}+20"

£(x - 50)(y - 20) = -100

"\Rightarrow xy-50y-20y+1000=-100
\\\Rightarrow xy-50y-20y=-1100"